# Black Holes and Wormholes Beyond Classical General Relativity

### v1

@@ Abstract
In the paper, only Static Spherically Symmetric space-times in four dimensions are considered within modified gravity models. The non-singular static metrics, including black holes not admitting a de Sitter core in the center and traversable wormholes, are reconsidered within a class of higher-order $F(R)$, satisfying the constraints $F(0)=\frac{dF}{dR}(0)=0$. Furthermore, by making use of the so-called effective field theory formulation of gravity, the quantum corrections to Einstein-Hilbert’s action due to higher-derivative terms related to curvature invariants are investigated. In particular, in the case of Einstein-Hilbert action plus cubic curvature Goroff-Sagnotti contribution, the second-order correction in the Goroff-Sagnotti coupling constant is computed. In general, it is shown that the effective metrics, namely Schwarzschild expression plus small quantum corrections, are related to black holes and not to traversable wormholes. In this framework, within the approximation considered, the resolution of singularity for $r=0$ is not accomplished. The related properties of these solutions are investigated.
preprint: APS/123-QED
@@ I Introduction
The quest for a consistent and predictive theory of quantum gravity remains one of the key challenges in modern physics [12, 13, 14]. Overcoming these challenges would represent significant progress towards a viable theory of quantum gravity. This will be capable of addressing unresolved issues such as the information paradox and the singularity problem that arises in classical and semi-classical approaches to gravitational phenomena. One intriguing proposal within the framework of superstring theory [15, 16] is the fuzzball proposal [17, 18], which offers a potential resolution to these issues. According to this proposal, a black hole (BH) is envisioned as a massive object composed of a vast number of microscopic strings, each with a minimal length extension on the order of the Planck scale. However, it is important to mention that the fuzzball proposal is not the only self-consistent and robust alternative for the generalization of BH physics incorporating quantum gravitational effects. Another important class of BHs can be derived within the framework of higher-derivative theories and non-local gravity [19]. These models address issues arising from the combination of quantum field theory and gravitation, such as non-renormalizability and non-unitarity, by incorporating higher-derivative terms in the Lagrangian of gravitational interaction. Interestingly, these additional terms can eliminate the undesirable features that impact the standard quantization of Einstein’s GR and may lead to potentially detectable effects in various physical phenomena.
@@
We note that BHs produce a ringdown waveform that is entirely determined by Quasinormal modes (QNMs), which are dependent solely on the mass and angular momentum of the BH. Therefore, the potential existence of additional “echoes” in the ringdown waveform could indicate the absence of an event horizon. This implies that horizonless compact objects (HCOs) such as gravastars  [24, 25, 26], boson stars  [27], or other exotic compact objects (ECOs)  [28, 24, 29, 30, 31, 32, 33, 34] cannot be ruled out as an alternative to BHs.
The paper is structured as follows: In Sec. II, we introduce non-singular BHs and wormholes (WH) in $F(R)$ modified gravity. Sec. III is dedicated to the WH solution in $R^{2}$ gravity, specifically a new spherically symmetric WH solution for $R^{2}$. In Sec. IV, making use of the effective field theory formulation of gravity, we will revisit the quantum gravitational corrections to the classical vacuum solution of GR. In Sec. V, we discuss the two-loop quantum-corrected BH related to the Goroff-Sagnotti contribution. Sec. VI covers $F(R)$ modified gravitational models. In Sec. V, we will explore the idea of a photon ring and shadow linked to a generic Static Spherically Symmetric (SSS) compact object, such as a BH or a WH. Finally, in Sec. VIII, we provide commentary on our results and offer perspectives on future research. In the paper we use the conventions $c=\frac{h}{2\pi}=1$, as well as the Newton constant $G_{N}=1$.
@@ II Non singular black holes and Wormholes in $F(R)$ modified gravity
In this section, we discuss nonsingular BH and WH solutions within a particular class of modified $F(R)$-gravity theories [35, 36]. We write down the equations of motion for $F(R)$- gravity [5, 37, 4, 6] in a vacuum, namely with vanishing stress tensor matter as,
$F_{R}(R)G_{\mu\nu}$ $=$ $\frac{1}{2}\left(F(R)-RF_{R}(R)\right)g_{\mu\nu}$ (1)
$+\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)F_{R}(R)\,.$
@@ II.1 A non-singular Black Hole
where $A(r)$ is a function of the radial coordinate $r$ only and $\ell$ is a positive length parameter, eventually on the Planckian size.
The associated Ricci scalar reads
@@
where the prime denotes a derivative with respect to the radial coordinate, $r$. Imposing $R=0$, one gets the second-order linear differential equation,
@@
$A(r)$ $=$ $1-\frac{C}{r}-\frac{3C\ell}{r^{5/2}}\left[\sqrt{r-\ell}\ln\left(\sqrt{r-\ell}+\sqrt{r}\right)-\sqrt{r}\right]$ (7)
$+\frac{Q\sqrt{r-\ell}}{r^{5/2}}\,,$
@@ III Wormhole solution in $R^{2}$ gravity
First, we recall the formalism we are going to use, see [42] for details.
A generic spherically symmetric space-time in isotropic coordinates reads
@@
On the horizon, the numerator is vanishing. In the case of a BH, the Kodama energy is non-vanishing, and $F(\rho)$ must have a double zero in $\rho$. In the case of WH, the Kodama energy is vanishing, and $F_{H}$ is different from zero.
As a check, we recall that the Schwarzschild solution in isotropic coordinates corresponds to
@@ III.1 A New Spherically Symmetric Wormhole Solution for $R^{2}$
$R=R_{1}+R_{2}\,,\quad R_{1}=\frac{2}{r}(1-B-rB^{\prime})\,,\quad$
$R_{2}=\frac{B}{A}\left(-A^{\prime\prime}+\frac{A^{\prime 2}}{2A}-\frac{A^{\prime}B^{\prime}}{2B}-\frac{2A^{\prime}}{r}\right).$ (15)
@@
For large $r$, one has $A=1-\frac{r_{s}}{(1+\lambda)r}+\mathcal{O}\left(\frac{r_{s}^{2}}{r^{2}}\right)$. As a consequence, we may put $2M=\frac{r_{s}}{1+\lambda}$. If $\lambda$ is very small, one has a WH which mimics the Schwarzschild BH. It should be noted that the quasi-local Misner -Sharp mass is $M_{MS}=\frac{r_{s}}{2}$.
Furthermore, it is easy to show that in isotropic coordinates, the solution (19) reads
$ds^{2}$ $=$ $-\frac{1}{(1+b_{1})^{2}}\left(b_{1}+\frac{4\rho-r_{s}}{4\rho+r_{s}}\right)^{2}dt^{2}$ (20)
$+\left(1+\frac{r_{s}}{4\rho}\right)^{4}\left(d\rho^{2}+\rho^{2}dS^{2}\right)\,.$
Thus, it does not belong to the Buchdahl class discussed in Appendix C.
A similar WH solution has been obtained in another gravitational model; see [43].
@@ IV Quantum gravitational corrections to classical Schwarzschild solution
The quantum gravitational adjustments to the emission of Hawking particles concept was first introduced by Calmet et al. [44, 45] using a Schwarzschild metric, akin to Hawking’s original work [46]. Expanding on this, they incorporated established quantum gravitational corrections to the metric and discovered that the quantum amplitude for the emission of a Hawking particle is influenced by quantum hair, which refers to the quantum gravitational correction to the classical Schwarzschild background. In this Section, we shall revisit the quantum gravitational corrections to the classical vacuum solution of GR.
@@ IV.1 The one-loop corrections
$F(\rho)$ $=$ $\left(1-\frac{m}{2\rho}+\frac{31m}{30\pi\rho^{3}}\right)^{2}\left(1+\frac{m}{2\rho}-\frac{31m}{30\pi\rho^{3}}\right)^{-2}\,,$
$\quad G(\rho)$ $=$ $\left(1+\frac{m}{2\rho}-\frac{7m}{30\pi\rho^{3}}\right)^{4}\,.$ (21)
Now, the horizon is shifted a little by the quantum correction and can be determined again by making use of the Eq. (10). As a result, one has
@@
Since $F(\rho_{H})>0$, namely, is not vanishing at the horizon, one is dealing with a traversable WH, and the new horizon becomes the WH throat. We recall that for static WH, the Hawking radiation is absent [42]. Furthermore, the presence of the throat implies that singularity in $r=0$ may be resolved.
Similar results have been recently obtained in [48, 49] within the so-called improved semi-classical approximation.
@@ IV.2 Effective Field Theory Extension of GR
Here, we consider the approach of [50], where an effective Lagrangian beyond GR has been investigated. We limit to the first correction, parametrized by dimensionless coupling constant $\varepsilon$. The effective Lagrangian is the Einstein-Hilbert plus a quadratic term in Curvature tensor. The complete equations of motion are reported in [50].
In the paper [51] a more general effective Lagrangian up to third order in the curvature has been investigated. We will come back to it in the next Section.
Within a metric given again by
@@
In the zero-order $\varepsilon=0$, one is dealing with GR, and the solution is unique and coincides with the Schwarzschild solution.
The solutions of the equations of motions up to the second order in $\varepsilon$ read [50]
@@
$f_{2}(r)$ $=$ $-\frac{1}{544}\left(\frac{2M}{r}\right)^{19}\bigg{(}198305-231709\frac{r}{M}$ (24b)
$+64584\frac{r^{2}}{M^{2}}\bigg{)}\,,$
@@
$g_{2}(r)$ $=$ $-\frac{1}{544}\left(\frac{2M}{r}\right)^{19}\bigg{(}2513399-2370222\frac{r}{M}$ (24d)
$+554472\frac{r^{2}}{M^{2}}\bigg{)}\,.$
@@
$r_{0}+\varepsilon r_{1}+\varepsilon^{2}r_{2}$ $=$ $2M-\varepsilon\left(\frac{2M}{r_{H}}\right)^{9}\left(\frac{67M}{4}-9r_{H}\right)$ (28)
$-\varepsilon^{2}r_{H}g_{2}(r_{H})\,.$
In order to check if one is dealing with a BH, one has to evaluate $f(r_{H})$. One has
@@
Making use of equations (24b) and (24d), one obtains
@@
As a consequence, $f(r_{H})=0$, and one is dealing again with a BH. In Appendix A, we will explore the first-order curvature coupling constant $\varepsilon$, which corresponds to the higher derivative model in the general form of the action.
@@ V Two-loop quantum corrected black hole
Since the one-loop corrections are vanishing in a vacuum, other and different quantum gravitational corrections to the Schwarzschild solution have been proposed [44, 45, 52]. In fact, as shown by Goroff and Sagnotti (GS) [53], the two-loop quantum correction is not vanishing in a vacuum. As a consequence, one may consider the following effective action consisting in the Einstein-Hilbert term plus the GS higher derivative counter term, an invariant which is cubic in the curvature tensor and not vanishing on shell see [44, 45, 52]. As already mentioned, a more complete analysis within the so-called low-energy effective theory of gravity approach is presented in [51]. Here, we limit ourselves to consider only the Goroff-Sagnotti term.
Let us begin again with the definition of the static metric form defined in Eq. (23). Then, the effective Lagrangian reads
@@
$\frac{1}{2\sqrt{fg}}\left[rg^{\prime}+g-1+\frac{\omega}{r^{4}}H^{3}\right]+\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{f}H-$
$\frac{d}{dr}\left[\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{f^{\prime}}H\right]+\frac{d^{2}}{dr^{2}}\left[\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{f^{\prime\prime}}H\right]=0\,,\quad$ (33a)
$\frac{-\sqrt{f}}{2\sqrt{g}g}\left[rg^{\prime}+g-1+\frac{\omega}{r^{4}}H^{3}\right]+\frac{\sqrt{f}}{\sqrt{g}}\left[1+3\omega\frac{H^{2}}{r^{4}}\partial_{g}H\right]$
$-\frac{d}{dr}\left[r\sqrt{\frac{f}{g}}+\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{g^{\prime}}H\right]=0\,.\quad$ (33b)
@@
$f(r)$ $=$ $f_{0}(r)+\omega f_{1}(r)+\omega^{2}f_{2}(r)+...\,,\quad$
$g(r)$ $=$ $g_{0}(r)+\omega g_{1}(r)+\omega^{2}g_{2}(r)+...\,.$ (34)
@@
with $M$ constant of integration.
Making use of Eq. (36), one can get the following set of equations
$\frac{r^{7}\left(rg_{1}^{\prime}+g_{1}\right)+6912M^{2}\left(2M-15(r-2M)\right)}{2r^{6}(r-2M)}\omega$
$+\mathcal{O}\left(\omega^{2}\right)=0\,,$ (37a)
$-\frac{r^{8}f_{1}^{\prime}+6912M^{2}(8M-3r)}{2r^{6}(r-2M)}\omega+\frac{2Mrf_{1}-r^{2}g_{1}}{2(r-2M)^{2}}\omega$
$+\mathcal{O}\left(\omega^{2}\right)=0\,.$ (37b)
@@
On the horizon $g(r_{H})=0$. This gives $r_{H}=2M+\omega r_{1}+..$, with $r_{1}=\frac{72}{M^{3}}$. As a result, $f(r_{H})=0$, and, to this first-order in $\omega$, one is dealing with a BH.
@@
$\frac{2^{16}3^{6}M^{3}\left(2^{2}683M^{2}-2319Mr+480r^{2}\right)}{r^{12}(2M-r)}\omega^{2}$
$-\frac{r\left(rg_{2}^{\prime}+g_{2}\right)}{2(2M-r)}\omega^{2}+\mathcal{O}\left(\omega^{3}\right)=0,$ (41a)
$-\left(2^{4}123M^{3}-2^{4}155M^{2}r+3^{2}115r^{2}M-2^{4}3^{2}r^{3}\right)\omega^{2}$
×$\times\frac{2^{16}3^{5}M^{3}}{r^{12}(r-2M)^{2}}-\frac{r^{2}\left((r-2M)f_{2}^{\prime}+g_{2}\right)-2Mrf_{2}}{2(r-2M)^{2}}\omega^{2}$
$+\mathcal{O}\left(\omega^{3}\right)=0\,.~{}\quad$ (41b)
@@
$g_{2}(r)$ $=$ $-\frac{2^{17}3^{5}M^{3}\left(7513M^{2}-3^{2}773Mr+2^{4}3^{2}11r^{2}\right)}{11r^{13}}\,,$
$f_{2}(r)=-\frac{2^{16}3^{5}M^{3}\left(2^{2}451M^{2}-3^{2}233Mr+576r^{2}\right)}{11r^{13}}\,.$
@@
$f_{2}(2M)=\frac{2^{4}\times 3^{5}\times 43}{11M^{8}}\,,~{}g_{2}(2M)=\frac{2^{4}\times 3^{5}\times 65}{11M^{8}}\,,$ (47)
@@
In the following, some remarks. First, the second-order corrected metric presented above is the main result of our paper. Up to the first order in $\omega$, our expressions for the metric are in agreement with several others works. First, Calmet et al [44, 45] have found
$f(r)$ $=$ $1-\frac{2M}{r}+a\frac{M^{3}}{r^{7}}\,,$
$g(r)$ $=$ $1-\frac{2M}{r}+b\frac{M^{2}}{r^{6}}\left(27-\frac{49M}{r}\right)\,,$ (50)
@@
$f(r)$ $=$ $1-\frac{r_{s}}{r}+2\lambda\frac{r_{s}^{2}}{r^{6}}\left(3-2\frac{r_{s}}{r}\right)\,,$
$g(r)$ $=$ $1-\frac{r_{s}}{r}+2\lambda\frac{r_{s}^{2}}{r^{6}}\left(9-8\frac{r_{s}}{r}\right)\,.$ (54)

### v2

@@ Abstract
Abstract: In the paper only Static Spherically Symmetric space-times in four dimension are considered within modified gravity models. The non-singular static metrics, including black holes not admitting a de Sitter core in the centre and traversable wormholes, are reconsidered within a class of higher-order $F(R)$, satisfying the constraints $F(0)=\frac{dF}{dR}(0)=0$. Furthermore, making use of the so called effective field theory formulation of gravity, the quantum corrections to Einstein-Hilbert action due to higher-derivative terms related to curvature invariants are investigated. In particular, in the case of Einstein-Hilbert action plus cubic curvature Goroff-Sagnotti contribution, the second order correction in the Goroff-Sagnotti coupling constant is computed. In general, it is shown that the effective metrics, namely Schwarschild expression plus small quantum corrections are related to black holes, and not to traversable wormholes. In this framework, within the approximation considered, the resolution of singularity for $r=0$ is not accomplished. The related properties of these solutions are investigated.
@@ I Introduction
The quest for a consistent and predictive theory of quantum gravity remains one of the key challenges in modern physics [14, 15]. Overcoming these challenges would represent significant progress towards a viable theory of quantum gravity. This will be capable of addressing unresolved issues such as the information paradox and the singularity problem that arises in classical and semi-classical approaches to gravitational phenomena. One intriguing proposal within the framework of superstring theory [16] is the fuzzball proposal [17, 18], which offers a potential resolution to these issues. According to this proposal, a black hole (BH) is envisioned as a massive object composed of a vast number of microscopic strings, each with a minimal length extension on the order of the Planck scale. However, it is important to mention that the fuzzball proposal is not the only self-consistent and robust alternative for the generalization of BH physics incorporating quantum gravitational effects. Another important class of BHs can be derived within the framework of higher-derivative theories and non-local gravity [19]. These models address issues arising from the combination of quantum field theory and gravitation such as non-renormalizability, non-unitarity by incorporating higher-derivative terms in the Lagrangian of gravitational interaction. Interestingly, these additional terms can eliminate the undesirable features that impact the standard quantization of Einstein’s GR and may lead to potentially detectable effects in various physical phenomena.
@@
We note that BHs produce a ringdown waveform that is entirely determined by Quasinormal modes (QNMs), which are dependent solely on the mass and angular momentum of the BH. Therefore, the potential existence of additional “echoes” in the ringdown waveform could indicate the absence of an event horizon. This implies that horizonless compact objects (HCOs) such as gravastars  [24, 25], boson stars  [26], or other exotic compact objects (ECOs)  [27, 28, 29, 30, 31, 32, 33] cannot be ruled out as an alternative to BHs. Lambiase et al. [34] examined how the quantum-gravity correction parameter $c_{6}$ at the third-order curvature affects black hole shadows and weak deflection angles. They achieved this by studying the photon sphere and using the finite-distance version of the Gauss-Bonnet theorem (GBT).
The paper is structured as follows: In Sec. II, we introduce non-singular BHs and wormholes (WH) in $F(R)$ modified gravity. Sec. III is dedicated to the WH solution in $R^{2}$ gravity, specifically a new spherically symmetric WH solution for $R^{2}$. In Sec. IV, making use of the effective field theory formulation of gravity, we will revisit the quantum gravitational corrections to the classical vacuum solution of GR. In Sec. V, we discuss the two-loop quantum-corrected BH related to the Goroff-Sagnotti contribution. Sec. VI covers $F(R)$ modified gravitational models. In Sec. VII, we will explore the idea of a photon ring and shadow linked to a generic Static Spherically Symmetric (SSS) compact object, such as a BH or a WH. Finally, in Sec. VIII, we provide commentary on our results and offer perspectives on future research. In the paper we use the conventions $c=\frac{h}{2\pi}=1$, as well as the Newton constant $G_{N}=1$.
@@ II Non singular black holes and Wormholes in $F(R)$ modified gravity
In this section, we discuss the nonsingular BH and WH solutions within a particular class of modified $F(R)$-gravity theories[35, 36]. We write down the equations of motion for $F(R)$- gravity [7, 39, 6, 40] in a vacuum, namely with vanishing stress tensor matter as,
$F_{R}(R)G_{\mu\nu}=\frac{1}{2}\left(F(R)-RF_{R}(R)\right)g_{\mu\nu}+\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)F_{R}(R)\,.$ (1)
@@
The Ricci tensor, $R_{\mu\nu}$ takes the form
$R_{\mu\nu}=\frac{\partial\Gamma^{\gamma}_{\mu\nu}}{\partial x^{\gamma}}-\frac{\partial\Gamma^{\gamma}_{\mu\gamma}}{\partial x^{\nu}}+\Gamma^{\lambda}_{\mu\nu}\Gamma^{\gamma}_{\lambda\gamma}-\Gamma^{\lambda}_{\gamma\mu}\Gamma^{\gamma}_{\nu\lambda}.$ (2)
The Christoffel symbol in the above equation is
$\Gamma^{\gamma}_{\mu\nu}=\frac{1}{2}g^{\gamma\sigma}\left(\frac{\partial g_{\sigma\nu}}{\partial x^{\mu}}+\frac{\partial g_{\sigma\mu}}{\partial x^{\nu}}-\frac{\partial g_{\mu\nu}}{\partial x^{\sigma}}\right),$ (3)
where the $R=g^{\mu\nu}R_{\mu\nu}$ is the Ricci scalar.
The Kretschmann scalar, $K=R^{\mu\nu\gamma\sigma}R_{\mu\nu\gamma\sigma}$ measures the curvature of spacetime near a black hole and is often used to analyze the nature of singularities. For a more detailed and rigorous understanding of this topic, one can refer to academic papers and research articles that delve into the mathematical and physical implications of the Kretschmann scalar in the context of singularities [41]. These papers offer in-depth discussions and clarity on the behavior of spacetime curvature at the singularity of a black hole.
@@ II.1 A non-singular Black Hole
where $A(r)$ is a function of the radial coordinate $r$ only and $\ell$ is a positive length parameter, eventually on the Planckian size. The associated Ricci scalar reads
@@
where the prime denotes a derivative with respect to the radial coordinate $r$. Imposing $R=0$, one gets the second-order linear differential equation as,
@@
$A(r)=1-\frac{C}{r}-\frac{3C\ell}{r^{5/2}}\left(\sqrt{r-\ell}\ln\left(\sqrt{r-\ell}+\sqrt{r}\right)-\sqrt{r}\right)+\frac{Q\sqrt{r-\ell}}{r^{5/2}}\,,$ (9)
@@
The regularity of the metric (6) can be understood following the remarks contained in [44]. The argument proceeds as follows. Introduce the new radial coordinate $\sigma$ defined by
$r^{2}=\sigma^{2}+\ell^{2}\,.$ (10)
Thus, one has
$ds^{2}=-A(r(\sigma))dt^{2}+\frac{d\sigma^{2}}{A(r(\sigma))}\left(1+\frac{\ell}{\sqrt{\sigma^{2}+\ell^{2}}}\right)+(\sigma^{2}+\ell^{2})dS^{2}\,.$ (11)
In this form, the regularity is manifest since $r=0$ is excluded as soon as $\ell>0$.
@@ III Wormhole solution in $R^{2}$ gravity
First, we recall the formalism we are going to use, see [45] for details. A generic spherically symmetric space-time in isotropic coordinates reads
@@
On the horizon, the numerator is vanishing. In the case of a BH, the Kodama energy is non-vanishing, and $F(\rho)$ must have a double zero in $\rho$. In the case of WH, the Kodama energy is vanishing, and $F_{H}$ is different from zero. As a check, we recall that the Schwarzschild solution in isotropic coordinates corresponds to
@@ III.1 A New Spherically Symmetric Wormhole Solution for $R^{2}$
$R=R_{1}+R_{2}\,,\quad R_{1}=\frac{2}{r}(1-B-rB^{\prime})\,,\quad R_{2}=\frac{B}{A}\left(-A^{\prime\prime}+\frac{A^{\prime 2}}{2A}-\frac{A^{\prime}B^{\prime}}{2B}-\frac{2A^{\prime}}{r}\right).$ (19)
@@
For large $r$, one has $A=1-\frac{r_{s}}{(1+\lambda)r}+\mathcal{O}\left(\frac{r_{s}^{2}}{r^{2}}\right)$. As a consequence, we may put $2M=\frac{r_{s}}{1+\lambda}$. If $\lambda$ is very small, one has a WH which mimics the Schwarzschild BH. It should be noted that the quasi-local Misner -Sharp mass is $M_{MS}=\frac{r_{s}}{2}$. Furthermore, it is easy to show that in isotropic coordinates, the solution (23) reads
$ds^{2}=-\frac{1}{(1+b_{1})^{2}}\left(b_{1}+\frac{4\rho-r_{s}}{4\rho+r_{s}}\right)^{2}dt^{2}+\left(1+\frac{r_{s}}{4\rho}\right)^{4}\left(d\rho^{2}+\rho^{2}dS^{2}\right)\,.$ (24)
Thus, it does not belong to the Buchdahl class discussed in Appendix C. A similar WH solution has been obtained in another gravitational model; Ref. [46].
@@ IV Quantum gravitational corrections to classical Schwarzschild solution
The quantum gravitational adjustments to the emission of Hawking particles concept was first introduced by Calmet et al. [47] using a Schwarzschild metric, akin to Hawking’s original work [48]. Expanding on this, they incorporated established quantum gravitational corrections to the metric and discovered that the quantum amplitude for the emission of a Hawking particle is influenced by quantum hair, which refers to the quantum gravitational correction to the classical Schwarzschild background. In this Section, we shall revisit the quantum gravitational corrections to the classical vacuum solution of GR.
@@ IV.1 The one-loop corrections
$F(\rho)=\left(1-\frac{m}{2\rho}+\frac{31m}{30\pi\rho^{3}}\right)^{2}\left(1+\frac{m}{2\rho}-\frac{31m}{30\pi\rho^{3}}\right)^{-2}\,,\quad G(\rho)=\left(1+\frac{m}{2\rho}-\frac{7m}{30\pi\rho^{3}}\right)^{4}\,.$ (25)
Now, the horizon is shifted a little by the quantum correction and can be determined again by making use of the equation (14). As a result, one has
@@
Since $F(\rho_{H})>0$, namely, is not vanishing at the horizon, one is dealing with a traversable WH, and the new horizon becomes the WH throat. We recall that for static WH, the Hawking radiation is absent [45]. Furthermore, the presence of the throat implies that singularity in $r=0$ may be resolved. Similar results have been recently obtained in [50, 51] within the so-called improved semi-classical approximation.
@@ IV.2 Effective Field Theory Extension of GR
Here, we consider the approach of [52], where an effective Lagrangian beyond GR has been investigated. We limit to the first correction, parametrized by dimensionless coupling constant $\varepsilon$. The effective Lagrangian is the Einstein-Hilbert plus a quadratic term in Curvature tensor. The complete equations of motion are reported in [52]. In Ref. [53] a more general effective Lagrangian up to third order in the curvature has been investigated. We will come back to it in the next section. Within a metric given again by
@@
in the zero-order $\varepsilon=0$, one is dealing with GR, and the solution is unique and coincides with the Schwarzschild solution. The solutions of the equations of motions up to the second order in $\varepsilon$ read [52]
@@
$f_{2}(r)=-\frac{1}{544}\left(\frac{2M}{r}\right)^{19}\left(198305-231709\frac{r}{M}+64584\frac{r^{2}}{M^{2}}\right)\,,$ (28b)
@@
$g_{2}(r)=-\frac{1}{544}\left(\frac{2M}{r}\right)^{19}\left(2513399-2370222\frac{r}{M}+554472\frac{r^{2}}{M^{2}}\right)\,.$ (28d)
Where $M$ be the mass of the black hole.
@@
$r_{0}+\varepsilon r_{1}+\varepsilon^{2}r_{2}=2M-\varepsilon\left(\frac{2M}{r_{H}}\right)^{9}\left(\frac{67M}{4}-9r_{H}\right)-\varepsilon^{2}r_{H}g_{2}(r_{H})\,.$ (32)
In order to check if one is dealing with a BH, then need to evaluate $f(r_{H})$. So,
@@
Making use of equations (28b) and (28d), one can obtain
@@
As a consequence, $f(r_{H})=0$, and one is dealing again with a BH.
@@ V Two-loop quantum corrected black hole
Since the one-loop corrections are vanishing in a vacuum, other and different quantum gravitational corrections to the Schwarzschild solution have been proposed [47, 54]. In fact, as shown by Goroff and Sagnotti (GS) [55], the two-loop quantum correction is not vanishing in a vacuum. As a consequence, one may consider the following effective action consisting in the Einstein-Hilbert term plus the GS higher derivative counter term, an invariant which is cubic in the curvature tensor and not vanishing on shell [Ref. [47, 54]]. As already mentioned, a more complete analysis within the so-called low-energy effective theory of gravity approach is presented in [53]. Here, we limit ourselves to consider only the Goroff-Sagnotti term. We begin with the definition of the static metric form defined in equation (27). Then the effective Lagrangian reads
@@
$\frac{1}{2\sqrt{fg}}\left(rg^{\prime}+g-1+\frac{\omega}{r^{4}}H^{3}\right)+\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{f}H-\frac{d}{dr}\left(\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{f^{\prime}}H\right)+\frac{d^{2}}{dr^{2}}\left(\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{f^{\prime\prime}}H\right)=0\,,$ (37a)
$-\frac{\sqrt{f}}{2\sqrt{g}g}\left(rg^{\prime}+g-1+\frac{\omega}{r^{4}}H^{3}\right)+\frac{\sqrt{f}}{\sqrt{g}}\left(1+3\omega\frac{H^{2}}{r^{4}}\partial_{g}H\right)-\frac{d}{dr}\left(r\sqrt{\frac{f}{g}}+\frac{3\omega\sqrt{f}H^{2}}{\sqrt{g}r^{4}}\partial_{g^{\prime}}H\right)=0\,.$ (37b)
@@
$f(r)=f_{0}(r)+\omega f_{1}(r)+\omega^{2}f_{2}(r)+...\,,\quad g(r)=g_{0}(r)+\omega g_{1}(r)+\omega^{2}g_{2}(r)+...\,.$ (38)
@@
with $M$ be the constant of integration. Making use of equation (40), one can get the following set of equations
$\frac{r^{7}\left(rg_{1}^{\prime}+g_{1}\right)+6912M^{2}\left(2M-15(r-2M)\right)}{2r^{6}(r-2M)}\omega+\mathcal{O}\left(\omega^{2}\right)=0\,,$ (41a)
$\frac{(2M-r)\left(r^{8}f_{1}^{\prime}+6912M^{2}(8M-3r)\right)+2Mr^{7}f_{1}+r^{8}(-g_{1})}{2r^{6}(r-2M)^{2}}\omega+\mathcal{O}\left(\omega^{2}\right)=0\,.$ (41b)
@@
On the horizon $g(r_{H})=0$. This gives $r_{H}=2M+\omega r_{1}+...$, with $r_{1}=\frac{72}{M^{3}}$. As a result, $f(r_{H})=0$, and, to this first-order in $\omega$, one is dealing with a BH.
@@
$\frac{2^{17}3^{6}M^{3}\left(2^{2}683M^{2}-2319Mr+480r^{2}\right)-r^{13}\left(rg_{2}^{\prime}+g_{2}\right)}{2r^{12}(2M-r)}\omega^{2}+\mathcal{O}\left(\omega^{3}\right)=0,$ (45a)
$-\frac{2^{17}3^{5}M^{3}\left(2^{4}123M^{3}-2^{4}155M^{2}r+3^{2}115r^{2}M-2^{4}3^{2}r^{3}\right)+[r^{14}\left((r-2M)f_{2}^{\prime}+g_{2}\right)-2Mr^{13}f_{2}]}{2r^{12}(r-2M)^{2}}\omega^{2}+\mathcal{O}\left(\omega^{3}\right)=0\,.~{}$
@@
$g_{2}(r)$ $=$ $-\frac{2^{17}3^{5}M^{3}\left(7513M^{2}-3^{2}773Mr+2^{4}3^{2}11r^{2}\right)}{11r^{13}}\,,$ (46a)
$f_{2}(r)=-\frac{2^{16}3^{5}M^{3}\left(2^{2}451M^{2}-3^{2}233Mr+576r^{2}\right)}{11r^{13}}\,.$ (46b)
@@
$f_{2}(2M)=\frac{2^{4}\times 3^{5}\times 43}{11M^{8}}\,,~{}~{}~{}g_{2}(2M)=\frac{2^{4}\times 3^{5}\times 65}{11M^{8}}\,,$ (51)
@@
Some remarks: First, the second-order corrected metric presented above is the main result of our paper. Up to the first order in $\omega$, our expressions for the metric are in agreement with several others works. First, Calmet et al [47] have found
$f(r)=1-\frac{2M}{r}+a\frac{M^{3}}{r^{7}}\,,\quad g(r)=1-\frac{2M}{r}+b\frac{M^{2}}{r^{6}}\left(27-\frac{49M}{r}\right)\,,$ (54)
@@
$f(r)=1-\frac{r_{s}}{r}+2\lambda\frac{r_{s}^{2}}{r^{6}}\left(3-2\frac{r_{s}}{r}\right)\,,\quad g(r)=1-\frac{r_{s}}{r}+2\lambda\frac{r_{s}^{2}}{r^{6}}\left(9-8\frac{r_{s}}{r}\right)\,.$ (58)